OpenStudy (anonymous):

what is 5 times 1 and 3/5

TheSmartOne (thesmartone):

Make 1 and 3/5 into an improper fraction :)

OpenStudy (arihangdu):

5.1(3/5)=3

OpenStudy (arihangdu):

5.1(3/5)=5(3/5)=15/5=3.